OpenStudy (anonymous):

on a grid, the distance between the points located at (1,-2) and (-9,3) is roughly equal to what?

OpenStudy (anonymous):

\[d = \sqrt {\left( {x_1 - x_2 } \right)^2 + \left( {y_1 - y_2 } \right)^2 }\] \[d = \sqrt {\left( {1 - (-9) } \right)^2 + \left( {-2 - 3 } \right)^2 }\] \[d = \sqrt {\left( {10 } \right)^2 + \left( {-5 } \right)^2 }\] \[d = \sqrt {100 + 25 }\] \[d = \sqrt {125 }\]

OpenStudy (anonymous):

can u simplify it from the last part?